OpenStudy (anonymous):

What is the discriminant of 3x^2 + 6x = 2? 12 18 42 60

OpenStudy (anonymous):

3x^2 + 6x = 2 -2 on both sides to get quadratic equation = 0 3x^2 + 6x -2 = 0 ax^2 + bx + c = 0 b^2 -4ac plug it in